Operations Report – September 5th

MDRS Complete Operations Report 05-SEPT-2021
Name of person filing report: Shannon Rupert
Reason for Report: Beginning of field season/work party daily update
Non-nominal systems: Old leak in the kitchen sink in Hab has resurfaced in the ground below the kitchen
Action taken for non-nominal systems: Nothing to report
Generator: Nothing to report
ScienceDome Dual Split: Operating nominally and I love having AC to escape to when it’s so hot in the afternoon
Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 76.5
Minimum 49
Maximum 100
Note on solar: work party crews are not power conserving, and the heat is adding to our power consumption
Diesel Reading – 0 %.
Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%
Ethanol Free Gasoline – not checked gallons
Water (loft tank): 50 gallons
Water Meter: n/a unit
Water (static tank) – 100 gallons
Static to Loft Pump used – yes
Water in GreenHab – 200 gallons, filled
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 200 gallons
Hab toilet tank emptied: no
Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge:
Currently charging: yes
Notes on rovers: Nothing to report
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). One of the 350’s is being used as a signReason for use: n/aOil Added? NoATV Fuel Used: Not checked Gals# Hours the ATVs were Used today: 0 Notes on ATVs: Nothing to report
HabCar used and why, where? Yes, hauling water
CrewCar used and why, where? Yes, the team took it to the Goblin Valley this afternoon and it was used at the storage units this morning. It will need to be cleaned before crew use.
Luna used and why, where? No
General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. We will work on radios tomorrow.
Campus-wide inspection, if action taken, what and why?

Volunteer vegetation around campus and MOOP have been removed.
I don’t know if I mentioned yesterday that we outlined in rocks a parking spot for visitors and borrowed the URC orange cones to stop people from coming up to MDRS.
Summary of general operations: We began work on the flooded storage units today but removing anything that has not been used in several years. We announced around town that it was first come, first served on anything outside. We will go tomorrow afternoon and empty more.

The goal is to have everything on shelves, then power wash the mud out, and hopefully have time to move all remaining materials into a single storage unit with nothing on the floor.

Summary of Hab operations: I remembered the leak in the kitchen sink today because the ground is wet again. That will need to be repaired before the first crew arrives.

Summary of Outpost operations: Septic and intern trailer blackwater tanks will be pumped on Thursday.

We finished the construction of the new bed in the intern trailer. It has a double bed on the bottom and a twin on top. They have the best view on Mars.

Summary of GreenHab operations: 60% shade cloth installed. The soil was added to the herb garden, the two greens containers, one of the carrot bins, plus 12 two-gallon pots and 21 one-gallon pots. We made two water runs for the GreenHab today.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Too hot. I am trying to get the team to embrace working in the early morning. So far, they aren’t embracing!

Questions, concerns, supplies needed and requests: Nothing at this time.

Shannon M Rupert, PhD
Director
Mars Desert Research Station
P.O. Box 82
Hanksville, Utah 84734

Supplemental Operations Report 04-SEPT-2021

MDRS Supplemental Operations Report 04-SEPT-2021

This is the second of two operations reports filed tonight.

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party

Non-nominal systems: This is a reminder of what systems needed repair when we left: Leak in Hab toilet, leak in upper deck water tank intake, GreenHab door, leak in ScienceDome under window with damage

Action taken for non-nominal systems: These are all issues that need to be addressed before the field season begins

Generator check, note if oil and coolant added: New generator has not been activated for autostart. I will do a test run on the machine in the new two weeks.

ScienceDome Dual Split: Set to automatically keep temperature at 77 degrees F/24 degrees C for the summer, and worked perfectly. I turned it lower today so we could escape in there when it got hot in the afternoon.

Solar— Nominal, providing all power

Diesel Reading – 0 %.

Propane Reading, main tank – 80 %

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72% (loss of 3%) over summer

Generator propane tank: 80%

Water (Outpost tank) – 200 gallons

Campus wide inspection, if action taken, what and why? Contacted contractor to pump septic tank at Outpost. We will make arrangements for the main septic to be pumped later this fall. We created a rock outlined parking lot at the main sign as the first step in our tourist design.

Volunteer vegetation around campus is currently being removed.

Summary of Hab operations: There has been significant damage to the exterior shell, which we will be painting this week. In some places it has cracked and I am hopeful the elastomeric paint will seal it again. There is also some evidence of water around the window of the upper deck. We will try to attach drip flashing after painting, and repair interior damage as soon as we can. Thee staircase is very rusted as it was never properly finished and sealed. That is something that should be done between crews. At minimum rust needs to be removed.

Summary of Outpost operations: We attacked the intern trailer with a vengeance this morning! The sofa and dining booth were removed, as well as all the window valences. The designed and purchased material to build a full-size bunk-bed for the front living area and a wall to divide the bedroom from the bathroom to allow more privacy when a group of people are living there. The dining booth will be replaced with a table and chairs. The walls will be painted and new insulated curtains put on the windows. Additionally, we need to make repairs to the water system and repair the main furnace. Finally, the trailer is listing several degrees on one side, so we need to level it.

Summary of GreenHab operations: Nothing to report except the potting soil and some plants were purchased.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Some paint touch up should be done and some leaks need to be sealed. It was a very humid summer as we had some rusting of materials in the building.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: It’s been hotter and getting hotter so we are keeping the dome cool, drinking a lot of water and making sure not to work in the sun during the hottest time of the day.

Questions, concerns, supplies needed and requests: We have requested an additional person or two from the URC to come to Hanksville and help with the clean out of the storage units. As you may know, Hanksville was hit with heavy rain and there was major flooding in a concentrated area in town. That area included the storage facility where MDRS and the URC rent two units. The entire area was flooded and our units have between 2 and ½ inches of wet mud completely covering them. A lot of the Ikea materials for the ScienceDome were on the floor and had varying degrees of mud and water damage. A complete inventory will be needed to determine what we have to replace. In the meantime, we need to completely clear out the units and power wash them this week and get everything we can either out or cleaned up and put away. The remodel of the Science Dome has been put on hold for now. The large trailer will be brought to the storage place so we can get rid of stuff.

Operations Report 4-Sept-2021

Work Party Crew Operations Report 04-Sept-2021

This is one of two operations reports filed today

SOL: N/A

Name of person filing report: Shannon Rupert

Non-nominal systems: Intern trailer water, intern trailer heat.

Notes on non-nominal systems: Working on repairs

Generator: Nothing to report

Hours run: N/A

Solar— SOC % Last 24 hours :

Average 85.2

Minimum 61

Maximum 100

Station Propane Reading – 80 %

Ethanol Free Gasoline – not checked gallons

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 200 gallons

Static to Loft Pump used – yes

Water in GreenHab –0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: Yes

Hours: 216.4

Beginning charge: 96

Ending charge: 97

Currently charging: yes

Notes: When we first arrived, Percy was parked outside my trailer and plugged in but the malfunction light was flashing and it was at 18% when it should have been at 100%. We couldn’t determine why it was not charging so we moved it to the RAM and tried to charge it again. It is slowly charging and the light has stopped blinking. Later we discovered that the 110 plug in the outpost power tower was tripped, probably due to heavy rain. Rover appears to be nominal now.

Notes on rovers: The 4 remaining rovers have had their batteries replaced and were supposed to be delivered this morning, but there was a lack of communication between parties and they did not get delivered. My assumption is that delivery will be rescheduled for sometime after the holiday weekend.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). The three 350’s. The Honda is still non-operational. The 300 is no longer at MDRS.

Reason for use: Test for battery and road worthiness.

Oil Added? No

ATV Fuel Used: Not checked Gals

# Hours the ATVs were Used today: 2

Notes on ATVs: The 350’s are all functional and in generally good operating order. We had to jump one, and tighten the battery cables on another but they all started. The tires were checked and air added as needed.

HabCar used and why, where? Not until this evening, when the new battery was installed. We will use it tomorrow to get water for the GreenHab,

CrewCar used and why, where? Yes, team took it to the Quarry and it was used to go check out the storage units in town.

Luna used and why, where? Yes, to Grand Junction for work party materials

General notes and comments: Grateful for all three vehicles today

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are as we left them in the station but without batteries, as the batteries were stored over summer in the ScienceDome and we have not put them back in the radios.

Summary of Hab operations: Mostly nominal. We have not turned on the hot water yet, but probably will tonight.

Summary of GreenHab operations: Potting soil was purchased We will add soil to the GreenHab. Fill the water tank and put up the shade cloth tomorrow. Seeds will be planted within the next couple of days.

Summary of ScienceDome operations: All nominal, no work done here.

Summary of RAM operations: This building really needs a good solid cleaning.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: We put out poison this summer to avoid rodent infestation, but most of it remained untouched. There was some evidence of pack rat visitors in the intern trailer and the Hab, but remaining buildings show no sign of rodents.

Questions, concerns and requests to Mission Support: Please add Marufa and Ashley to this eList. They are our Operations Interns for this season and they have been copied on this email.

We have requested additional personal from the URC team to assist with the cleaning and mud removal in the storage units.

Crew 235 – Final Mission Summary

MDRS Crew 235 Spaceward Bound Mission Summary

May 7, 2021

Summary Title: Generally Speaking, this is how it went

Authors’ names: Thomas Quayle (Science Lion Simba) and Jen Carver-Hunter

Sol Activity Summary:

At the start of our mission, nine strangers came together with different hopes and expectations, all with the same end goal – come together with other like-minded individuals within the education community and emerge with new skills or ideas. Amazingly, every step of our journey together built the outcome that would meet and exceed everyone’s expectations.

Many of the crew interactions and experiences have been recorded in the journal reports that have been submitted each day of the mission. This report will focus on the accomplishment of our mission goals, which included completing a microbiology experiment and geologic fieldwork for collection, comparison, and analysis.

For our microbiology experiment, crew prepared agar, which they used to grow the fungi and bacteria that were swabbed from different areas inside the Hab. After allowing the specimens to incubate for 3 days, crew sorted and categorized the agar to identify the locations that were swabbed. This was done in a blind categorization so both crew teams approached it through flexible problem solving. The first team used a matrix system for identification while the second team used process of elimination and benchmarks to identify the tested locations for each sample. Every crew member felt the microbiology experiment was a success in that it would be transferable to a K-12 classroom environment.

For our geologic fieldwork experience, crew members used a presentation about Mars geology by Dr Rupert as a spring board for identification of MDRS rock samples that represent the geology of Mars. We then used 3 EVAs over the next 3 days to collect rock samples we will use to create our MDRS-Mars rock collections. On the first EVAs, crew members traveled to Kissing Camels Ridge and followed the ridge to the west to collect petrified wood, which we were using as an analogous sample for the sulphur located on Mars. Crew members were delighted to find a diverse variety of samples, including a wide range of colors and textures. On the second EVAs, crew members traveled to Barainca Butte to collect lava, basalt, and granite samples. We were also able to collect gryphaea samples during this EVA as an added bonus! The second stop for this EVA was near Zubrin’s head to collect sandstone blueberries from a blueberry nursery. This geologic phenomenon was one that continued to delight and amaze crew members. On the third set of EVAs, crew members traveled to Copernicus Overlook to see the big picture of MDRS geology. After the overlook, crews returned east on Brahe Hwy, looking for the glint of gypsum near Beige Moon. Every crew member was surprised and excited to see the glinting shine that covered the ground since it was not observable on the way to the overlook. Once gypsum was collected, crews continued on to Cowboy Corner to collect the sandstone samples that would complete their collections.

All crew members agree that each facet of our experience at MDRS contributed to a successful mission. In addition to completing our science goals, we were able to complete all of our Sim goals as well as our collaboration/networking and teambuilding goals. We are grateful for the opportunity to learn and grow in this once-in-a-lifetime experience. Our lives are changed for the better, and we are all eager to return to our classrooms to share our week of Sim on Mars. In addition, we will continue working with Dr Rupert to develop curriculum and materials that will allow MDRS to reach teachers far beyond the scope of the Spaceward Bound Utah program.

 

Operations Report – May 6th

Crew 235 Operations Report 6-MAY-2021

SOL: 4

Name of person filing report: Atila Meszaros

Non-nominal systems: ATV-2

Notes on non-nominal systems: ATV-2 gas pedal system is broken. Once you press it, it would get stuck, instead of retracting, and it could lead to an accident. There’s also a shift problem, but we suspect is related to the gas pedal system interfering with the shift change. As we leave on Saturday, the ATVs will have to be tuned up before next season.

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 92% at 20:09 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 70% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 151161.3 units

Water (static tank): 250 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 211.9

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#4 and EVA#5

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2 hours

Notes on ATVs: ATV 2 gs pedal out of service

HabCar used and why, where?: Yes, to go to town for supplies.

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #5.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: General cleaning and bacterial disposal.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Battery replaced and now rotation system is working.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight, Bernard!

EVA Report – May 6th

Crew 235 EVA Report 06-May-2021

EVA # 6, 7

Author: Jen Carver-Hunter

Purpose of EVA: Continuation of rock sample collection

EVA 6 Start time: 0800
EVA 6 End time: 0900
EVA 7 Start Time: 0930
EVA 7 End time: 1030

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimens were gypsum and sandstone. The first planned stop for the EVA was the overlook at Copernicus Hwy, and the second planned stop on return to the Hab was Cowboy Corner.

We also decided to use these EVAs as an opportunity for the new crew to conduct an EVA without an MDRS-experienced leader. For this reason, we included a map lesson refresher prior to the EVA to ensure that all crew members would be familiar with the geologic features.

The first EVA crew (EVA 6) left the Hab in vehicles at 0800 and followed Cow Dung Rd north to Brahe Hwy. They missed the turn to Brahe Hwy and found themselves at the quarry, where they turned around and made their way to Brahe Hwy.

Then they followed Brahe Hwy to Copernicus Hwy where they stopped for a chance for a 360-degree look at the formations in the area. Then they reboarded vehicles and returned east on Brahe Hwy toward Cow Dung Road until they spotted the glint of the sun shining off the gypsum on the ground.

They collected samples of gypsum near Beige Moon before continuing on to Cowboy Corner, where they decided to not stop for sandstone collection. They also needed to release a packrat near the reservoir. After releasing it, they returned to vehicles and came back to the Hab without incident.

The second EVA crew (EVA 7) left the Hab in vehicles at 0930 and also followed Cow Dung Road north to Brahe Hwy. They did not miss the turn, so they executed a perfect left turn toward Copernicus Hwy, and stopped at the intersection of the two roads to take in the sights at the overlook.

The crew then started up their vehicles (one rover and one ATV) to start their trip back to the Hab. They also stopped near Beige Moon to collect gypsum samples. They elected to also stop at Cowboy Corner to collect sandstone samples before returning to the Hab without incident.

Destination: The overlook at Copernicus Hwy/Brahe Hwy; Beige Moon, Cowboy Corner

Coordinates (use UTM NAD27 CONUS): Destination 1 – 516100, 4254400; Destination 2 – 516500, 4254400; Destination 3 – 518350, 4252700

EVA 4 Participants: Jeff, Randall, Thomas

EVA 5 Participants: Krysta, Jennifer, Allison

Road(s) and routes per MDRS Map: Entrance Rd to Cow Dung Rd to Brahe Hwy

Mode of travel: Pathfinder Rover and 2 ATVs (only 1 ATV used on the second EVA)

Journalist Report – May 6th

Crew 235 Journalist Report 06May2021
Author: Jeff Streba, Crew Journalist
Sol 4: Log Entry

Another beautiful day on Mars.
We had an early morning EVA out to the overlook and then we collected Gypsum and Sandstone samples. Because of the lack of functional suits, only 3 people went on
each EVA.

While one EVA team was out the other team analyzed bacteria. We were all impressed with the results.

When everyone was together, we discussed how to incorporate both the bacteria experiment and rockhounding into our classrooms.

We then ended our sim

Journalist Report – May 5th

Crew 235 Journalist Report 05May2021
Authors: Thomas Quayle and Randall Gibson, Crew Journalist

Sol 3: Generally Speaking, It Goes Like This

Morning began with a brisk wind and another species of Rodentia captured. The crew has been remarkably passive about the recurring nature of the event. What began as a mixed curiosity has evolved into a morning routine, complete with EVA procedures, and smiles.

However, the first order of the day: coffee and another round of lively discussion surrounded the breakfast table. Today’s topic was the politics of Earth. We are hopeful they will come to solve their problems in time. The morning fair consisted of sausage and eggs. It has been great to enjoy the variety of meals that each of the crew members has been willing to jump in and prepare.

Little did the crew suspect that the quiet morning would not remain so for long. Even the most well-planned EVA cannot account for everything a crew can prepare for or expect. Their mission was to search for specific geologic samples in a region that was located nearly twice as far as the previous day’s activities.

More than an hour had passed from when the first EVA team had departed when a call came over the radio, “Hab, Hab!..”. The other half of the crew at the Hab quickly responded and was informed that the EVA was being aborted. Within minutes of the radio communique, the EVA team had arrived.

Thankfully no one had been injured, but a malfunction in the power pack of the Commander’s spacesuit (redacted) had occurred that seriously affected the ability for air to be recirculated. While it is not yet known where the malfunction took place or how it was missed, the event has impacted our EVA assessments for the duration of our mission. Each team will now have to continue with one crew member volunteering to abstain from participation. In the end, the second EVA team was able to successfully manage the completion of objectives scheduled for the day.

After the second EVA returned to the Hab, the crew reconvened as a single group to rest and clean the remains of lunch until the next event of the day. Shannon led a crew discussion on the assembly of Estes Rockets and the pitfalls to consider when introducing the topic to students. Each member of the crew was able to construct their own prototype, which should be available for launch tomorrow.

The crew then joined Shannon in the Science Dome to compare the samples that had been collected across multiple EVA’s. In the last significant event of the day, the astronomy box was retrieved from the outpost. This was the final component required to activate the solar observatory, which the crew is looking forward to using tomorrow.

As a social component, the crew finalized a Lego build that had been started the previous day. The genders and position of the Lego figures were debated and rearranged by the crew. The dinner discussion covered several topics, including the similarities to the Lewis and Clark expedition.

Jennifer prepared an incredible culinary masterpiece of southern fried chicken, fried potatoes, biscuits, and gravy by using dehydrated ingredients. This was surely the best meal yet to be prepared on this planet, and she shall be henceforth be known as the “Mars”tha Stewart.

Sol Summary – May 5th

Crew 235 Sol Summary Report 05-May-2021

Sol: 3

Summary Title: Rocks and Rockets!

Author’s name: Atila Meszaros

Mission Status: Crew Biologist

Sol Activity Summary:

We continued or geology hunting as EVA team #1 and #2 went South and reached the 1101 road to collect granite, lava in its different color variation, asphalt, devil’s toenail, and more to be analyzed by our expert, Dr. Shannon Rupert.

The first EVA encountered a small incident (check the EVA report for full explanation), but we are very close to finishing our geology hunt and gathering.

After lunch, we decided to start building our rockets and launching system for tomorrow’s project! Jennifer and Krysta are making dinner as we prepare the reports and get ready for a great night trying new activities and games.

Look Ahead Plan: Rock hunting EVAs in the morning, rocket launching, and going out of sim!

Anomalies in work: None.

Weather: Fully sunny, but with a great temperature throughout the day.

Crew Physical Status: Healthy. Finally learned how to work with dehydrated food and happy with the great meals and bread.

EVA: Tomorrow EVAs #5 and #6

Reports to be filed: Operations Report, EVA Report, EVA Request, Crew Journalist

Support Requested: Nothing to report.